### v2

@@ 6 Structural Pre-merger Analysis
In this section, we present a merger simulation framework that incorporates empirical insights from our difference-in-differences estimations in Sections 3 and 5. This approach connects our reduced-form findings with structural modeling to better understand the price and welfare implications when two airlines merge. Combining these complementary methodologies, we aim to address some of the limitations inherent in either approach used in isolation.
We begin by estimating a discrete choice model of airline demand and costs using pre-merger data and then use the estimates to determine post-merger prices. Then we use DiD estimated price effects (Table 3 Panel B) to calibrate efficiency gains and incorporate estimated changes in competitive conduct (as implied by our analysis in Section 4), we create a more comprehensive framework for merger evaluation. We note that the DiD analysis examines average market prices, but the structural approach considers firm-specific pricing decisions. We will return to this later when discussing the counterfactual analysis.
@@ 6.1 Demand Specification
We divide the choice set in each market $m\times t$ into two primary groups: the outside option ($g=0$), which corresponds to not flying, and inside options ($g=1$), which corresponds to flying using one of the flight services offered by the airlines. We further divide the inside options into nonstop services ($h=1$) and connecting services ($h=2$). Let $J_{hgmt}$ represent the set of airline offerings in subgroup $h$ of a group $g$ in market $m\times t$.
@@
where $X_{jmt}$ represents a vector of product characteristics, $p_{jmt}$ denotes price, and (${\boldsymbol{\phi}},\varrho$) are preference parameters, $\xi_{j}$ are carrier specific indicator variables, $\xi_{t}$ are year-quarter indicator variables, and $\xi_{jmt}$ captures product characteristics unobserved by the econometrician. Crucially, we maintain that $\nu_{ijmt}$ is the unobserved utility component corresponding to flying with one airline and is given by $\nu_{ijmt}=\varepsilon_{igmt}+(1-\sigma_{2})\varepsilon_{ihgmt}+(1-\sigma_{1})\varepsilon_{ijmt},$ where $\varepsilon_{igm}$, $\varepsilon_{ihgm}$, and $\varepsilon_{ijm}$ are independent random variables that follow extreme value distribution, consistent with Berry (1994) and Verboven (1996).
@@
Following Björnerstedt and Verboven (2014, 2016), we can write the market share (choice probability) for an airline service $j\in J_{hgmt}$ as:
@@ 6.2 Estimation Results
Table 7 presents the estimated own- and cross-price elasticities and nesting parameters across the three major airline mergers. Our sample includes all pairwise combinations of the top 50 populated origin and destination markets, restricted to three years leading up to the merger. Our estimated elasticities reveal significant market power and clear substitution patterns. Own-price elasticities ($M_{ejj}$) range from -4.07 to -4.57, indicating that a 1% increase in price leads to approximately a $4-4.5\%$ decrease in demand. The Delta-Northwest merger period exhibits the highest own-price elasticity (-4.57). These estimates align with previous studies: Berry and Jia (2010) find the own price elasticity for leisure passengers is $-5.01$, Ciliberto et al. (2021) estimate the median price elasticity to be $[-4.03,-3.86]$, Bontemps et al. (2023) estimate it at $-3.78$, and Aryal et al. (2024) estimate it to be $-3.90$. The nested structure shows the strongest substitution between similar service types (elasticities 0.26-0.97), weaker substitution across service types (0.01-0.11), and minimal substitution with the outside option (0.002-0.005).
@@ 6.3 Merger Simulations
We conduct merger simulations with the estimated demand and cost parameters by removing one airline and analyzing the resulting price effects. While such simulations would be straightforward for homogeneous goods with symmetric firms, the airline industry presents additional complexity due to product differentiation and heterogeneous costs that may change post-merger through efficiency gains.
Even maintaining our assumptions of exogenous market structure and merger decisions, simulating post-merger outcomes requires careful consideration of demand characteristics and cost structures. We address this challenge through multiple counterfactual scenarios. For product offerings, we begin with a baseline case where the merged entity maintains both pre-merger product lines, then examine scenarios where it retains only one brand. For costs, following Ciliberto et al. (2021), we explore three scenarios: maintaining pre-merger costs, adopting the minimum costs between merging carriers (i.e., the maximum efficiency), and using the average of pre-merger costs. Additionally, we leverage our DiD estimates to discipline the implied efficiency gains, allowing us to trace price effects across various levels of market competition by varying conduct parameters.
@@
  • Notes: This table presents counterfactual merger analysis based on estimated discrete choice demand parameters and recovered pre-merger marginal costs. We simulate post-merger equilibrium prices (in hundreds of USD) for each merger under various cost and market structure scenarios. The cost scenarios consider three possibilities: the merging entities keep their pre-merger costs, which is relevant only when both merging firms remain in the market; the merged firm adopts the smallest of the merging firms’ costs; or the merged firm operates with the average of the merging firms’ pre-merger costs.
@@
Suppose the airlines can reach the most optimistic efficiency scenario. In that case, it presents a puzzle: How do we reconcile the prediction of lower prices here with the earlier findings of stable post-merger prices? As suggested earlier, one possible explanation is that the firm’s strategic conduct changed after the merger.
@@
Note: These contour plots show predicted post-merger prices (in hundreds of USD) as a function of pre- and post-merger conduct parameters. The left panel assumes merged firms achieve maximum cost efficiency by adopting the lower marginal costs of the merging parties. The right panel additionally allows the merged firm to inherit the more favorable demand characteristics. The darkness of the shades indicates price levels. Conduct ranges from perfect competition (0) to full collusion (1), where pre-merger conduct is on the x-axis, and post-merger conduct is on the y-axis.
Figure 3 illustrates this tradeoff through contour plots of predicted price (in hundreds of US dollars) effects. We compare two scenarios: one where firms achieve maximum cost efficiency (left panel) and another where, in addition to the maximum efficiency, the merged firm inherits the superior demand characteristics of the two merging firms (right panel). While cost efficiencies can drive price reductions, changes in post-merger conduct often dominate these benefits. Demand-side improvements are relatively minor compared to cost efficiencies and conduct changes. Lastly, we note that the simulated maximum prices of approximately \$229 in Figure 3 align closely with the observed average prices in treated markets (see Table 2), suggesting the model generates economically meaningful predictions. The contour plots consistently indicate higher post-merger conduct across both efficiency scenarios, though the magnitude of price effects varies with the assumed efficiency gains.
This empirical exercise reveals a fundamental tension in evaluating the three airline mergers. The estimated price stability could reflect either large efficiency gains offset by increased market power or modest but opposing efficiency and competitive conduct changes. This aspect creates a dilemma for merging airlines - claiming significant efficiencies implies accepting evidence of increased market power while arguing for minimal competitive effects undermines the efficiency justification for the merger. A unified approach makes this tradeoff explicit.
@@ 7 Conclusion
Looking ahead, promising directions for future research include extending our framework to incorporate endogenous market structure (e.g., Ciliberto et al., 2021; Fan and Yang, Forthcoming), product characteristics and allowing these to change with mergers (e.g., Sweeting, 2010; Fan, 2013; Wollmann, 2018; Li et al., 2022; Bontemps et al., 2023; Atalay et al., 2024). The challenge will be maintaining our current approach’s transparency while adding these additional features. Our methodology of connecting reduced-form and structural approaches provides a valuable template for such extensions.
@@ A.3 Additional Data Tables
  • Notes: Summary statistics table of the variables (and their units of measurement) used to estimate the structural model in Section 4. For each merger, the sample includes all pairwise combination of the top 50 populated origin and destination markets, restricted to three years leading up to the said merger.
@@
  • Notes: Summary statistics table of the variables used to estimate the nested-logit discrete choice model in Section 6, separated by the merger. For each merger, the sample includes all pairwise combination of the top 50 populated origin and destination markets, restricted to three years leading up to the said merger.
@@ B.1 DiD Estimation with Levels
  • Notes: This table reports DiD estimates from Equation (1) for airfare, passenger volume and offered seats. These variables are measured in levels, where seats are in 1,000. The analysis is structured such that the first three columns show the estimates for the three mergers separately, and the last column shows the estimates for the “pooled” sample. Panel A presents results using mean-deviation fixed effects, and Panel B presents results using the first differences method. ${}^{***}:p<0.01$; ${}^{**}:p<0.05$.
The estimates are displayed in Table B.1. Panel A shows substantial merger effects when using the fixed effects approach. The DL-NW merger reduced fares by \$6.69 ($\approx 3\%=(6.69/225.39)\%$ relative to pre-merger fares) while increasing passengers by 3,728 and capacity by 34,000 seats per market. These effects are both economically and statistically significant. The UA-CO merger showed minimal price effects (\$1.76 increase, statistically insignificant) but similar capacity expansion. The AA-US merger generated the largest fare reduction of \$16.14 (6.8%) with a significant increase in passenger volume and offered seats.
However, these effects largely disappear under the first differences (Panel B) method. The DL-NW fare effect drops to an insignificant -\$0.77, while UA-CO shows a modest -\$1.66 decrease. The AA-US fare effect shrinks to an insignificant -\$1.30 (from a substantial increase of \$32.9 under fixed effects). Similarly, passenger and capacity effects attenuate by 80-85%, though remaining statistically significant.
These findings provide evidence of stable post-merger prices and align with our main findings in Section 3 Table 3, indicating more modest merger effects than traditional fixed effects approaches might suggest.
@@ B.2 Analysis at the Airline-Market Level
For our main analysis presented in Table 3, we estimated Equation (1) at the market level, where each observation represents a unique market-time pair (e.g., the logarithm of average flight prices for Atlanta to St. Louis market at a given time). As a robustness check, we consider the fixed effects (mean deviation) estimator to examine whether the competitive merger effects documented in Table 3 Panel A might have been driven by market-level aggregation. In particular, we extend our analysis by estimating merger effects at the more granular market-airline level. In this specification, each observation corresponds to a unique market-airline-time combination—for instance, the logarithm of average prices for flights operated by Delta in the ATL-STL market at a specific period.
Our primary variable of interest remains $(\text{Treated}_{mt}\times\text{Post-Merger}_{mt})$ as defined in Equation (1). However, instead of incorporating market fixed effects, we include “market-airline” fixed effects while retaining the time fixed effects. Similarly, the % nonstop now measures the percentage of nonstop passengers for a specific airline in a given market.
@@ B.3 Pre- and Post-Merger Windows
  • Notes: This table reports DiD estimates from Equation (1) for airfare, passenger volume and offered seats. Panel A presents results using one month before/after as relevant period to estimate merger effects across, and Panel B presents results using two month before/after as relevant period. ${}^{***}:p<0.01$; ${}^{**}:p<0.05$.
We present the results in Table B.3. Panel A shows that when using only one month before/after as the relevant period, the merger effects on airfares are imprecisely estimated across all merger cases. The effects on passenger volume appear mixed, with significant negative effects for DL-NW, insignificant effects for UA-CO, and marginally significant positive effects for AA-US. However, capacity (offered seats) consistently shows significant increases in overlap routes across all mergers. Panel B presents results using a two-month window. The pattern is similar—airfare effects remain statistically insignificant, passenger volume effects vary across merger cases, and capacity effects maintain positive significance across all mergers. These findings suggest that the immediate effects of mergers differ from longer-term impacts, with capacity adjustments occurring more rapidly than price and demand responses. The consistent positive effect on offered seats across both time windows indicates that merged airlines quickly increased capacity in overlapping markets. However, the corresponding price adjustments and passenger volumes may require more time to materialize or stabilize.
@@ B.5 Pre-trend Analysis
In Figure B.1, we present the estimates from pre-trend analysis for 2 to 1 markets, mirroring the methodology employed in Figure 2 for 3 to 2 markets, except for UA-CO merger, there are not any 2 to 1 treated markets. The results reveal that the merger coefficients for price and passenger volume exhibit substantial instability in pre-merger periods across the two mergers. This instability suggests a violation of the parallel trends assumption.
@@ B.6 Synthetic DiD
Recent methodological innovations offer a promising avenue to address this issue. The synthetic control method (Abadie and Gardeazabal, 2003) extended to the DiD framework by Arkhangelsky et al. (2021) provides a data-driven approach to constructing control markets. This method systematically combines potential control markets to create a synthetic comparison group that better approximates the counterfactual of the treated markets. Providing a systematic method to construct control groups reduces researcher discretion in market selection and relaxes the rigid assumptions about trend specifications inherent in our previous analysis. For instance, when analyzing the DL-NW merger, a treated market with pre-merger characteristics similar to DL-NW overlap markets would receive a higher weight than a dissimilar market without parametric assumptions about the evolution of market outcomes.
@@
We apply this approach using the implementation package developed by Clarke et al. (2023) to reassess the findings in Table 3. The estimates are presented in Table B.5. These effects suggest three key results. First, we find limited evidence of merger effects on prices and passenger volumes. For instance, the UA-CO merger shows a minimal fare effect of -0.4% in levels and -0.8% in first differences (the latter being marginally significant). The passenger volume effects are similarly modest, approximately 1% in both specifications and statistically indistinguishable from zero. These estimates suggest economically negligible price impacts from the merger.

### v3

@@ 6 Structural Pre-merger Analysis
In this section, we present a merger simulation framework that incorporates empirical insights from our difference-in-differences estimations in Sections 3 and 5. This approach connects our reduced-form findings with structural modeling to better understand the price and welfare implications when two airlines merge. Combining these complementary methodologies, we aim to address some limitations inherent in either approach used in isolation.
We begin by estimating a discrete choice model of airline demand and costs using pre-merger data and then use the estimates to determine post-merger prices. Then, by treating the DiD estimates (Table 3, Panel B) as constraints that these simulated post-merger prices must rationalize through combinations of efficiency gains and competitive conduct changes (as analyzed in Section 4), we create a more comprehensive framework for merger evaluation. We note that the DiD analysis examines average market prices, but the structural approach considers firm-specific pricing decisions. We will return to this later when discussing the counterfactual analysis.
@@ 6.1 Demand Specification
We divide the choice set in each market $m\times t$ into two primary groups: the outside option ($g=0$), which corresponds to not flying, and the inside option ($g=1$), which corresponds to using one of the flight services offered by the airlines. We further divide the inside options into nonstop services ($h=1$) and connecting services ($h=2$). Let $J_{hgmt}$ represent the set of airline offerings in subgroup $h$ of a group $g$ in market $m\times t$.
@@
where $X_{jmt}$ represents a vector of product characteristics, $p_{jmt}$ denotes price, and (${\boldsymbol{\phi}},\varrho$) are preference parameters, $\xi_{j}$ are carrier-specific indicator variables, $\xi_{t}$ are year-quarter indicator variables, and $\xi_{jmt}$ captures product characteristics unobserved by the econometrician. Crucially, we maintain that $\nu_{ijmt}$ is the unobserved utility component corresponding to flying with one airline and is given by $\nu_{ijmt}=\varepsilon_{igmt}+(1-\sigma_{2})\varepsilon_{ihgmt}+(1-\sigma_{1})\varepsilon_{ijmt},$ where $\varepsilon_{igm}$, $\varepsilon_{ihgm}$, and $\varepsilon_{ijm}$ are independent random variables that follow extreme value distribution, consistent with Berry (1994) and Verboven (1996).
@@
Following Björnerstedt and Verboven (2014, 2016), we can write the market share (choice probability) for an airline service $j\in J_{hgmt}$ as
@@ 6.2 Estimation Results
Table 7 presents the estimated own- and cross-price elasticities and nesting parameters across the three major airline mergers. Our sample includes all pairwise combinations of the top 50 populated origin and destination markets, restricted to three years leading up to the merger. Our estimated elasticities reveal significant market power and clear substitution patterns. Own-price elasticities ($M_{ejj}$) range from -4.07 to -4.57, indicating that a 1% increase in price leads to approximately a $4-4.5\%$ decrease in demand. The Delta-Northwest merger period exhibits the highest own-price elasticity (-4.57). These estimates align with previous studies: Berry and Jia (2010) find the own-price elasticity for leisure passengers is $-5.01$, Ciliberto et al. (2021) estimate the median price elasticity to be $[-4.03,-3.86]$, Bontemps et al. (2023) estimate it at $-3.78$, and Aryal et al. (2024) estimate it to be $-3.90$. The nested structure shows the strongest substitution between similar service types (elasticities 0.26-0.97), weaker substitution across service types (0.01-0.11), and minimal substitution with the outside option (0.002-0.005).
@@ 6.3 Merger Simulations
We conduct merger simulations with the estimated demand and cost parameters by removing one airline and analyzing the resulting price effects. While such simulations would be straightforward for homogeneous goods with symmetric firms, the airline industry presents additional complexity due to product differentiation and heterogeneous costs that may change after the merger.
Even under our assumptions of exogenous market structure and merger decisions, simulating post-merger outcomes requires that we carefully consider demand characteristics and cost structures. We address this challenge through multiple counterfactual scenarios. For product offerings, we begin with a baseline case where the merged entity maintains both pre-merger product lines, then examine scenarios where it retains only one brand. For costs, following Ciliberto et al. (2021), we explore three scenarios: maintaining pre-merger costs, adopting the minimum costs between merging carriers (henceforth, the maximum efficiency), and using the average of pre-merger costs. Also, using the DiD estimates with the first difference as benchmarks, we explore cost and competitive conduct changes to reconcile the predicted post-merger prices and the DiD estimates.
@@
  • Notes: This table presents counterfactual post-merger prices (expressed in hundreds of USD) for each merger based on estimated discrete choice demand parameters and recovered pre-merger marginal costs and various cost and market structure scenarios. The cost scenarios consider three possibilities: the merging entities keep their pre-merger costs, which is relevant only when both merging firms remain in the market; the merged firm adopts the smallest of the merging firms’ costs; or the merged firm operates with the average of the merging firms’ pre-merger costs.
@@
Suppose the airlines can reach the most optimistic efficiency scenario. In that case, it presents a puzzle: How do we reconcile the prediction of lower prices here with the earlier findings of stable post-merger prices? As suggested earlier, the firm’s strategic conduct may have changed after the merger.
@@
Note: These contour plots show predicted post-merger prices (in hundreds of USD) as a function of pre- and post-merger conduct parameters. The left panel assumes merged firms achieve maximum cost efficiency by adopting the lower marginal costs of the merging parties. The right panel also allows the merged firm to inherit the more favorable demand characteristics. The darkness of the shades indicates price levels. Conduct ranges from perfect competition (0) to full collusion (1), where pre-merger conduct is on the x-axis, and post-merger conduct is on the y-axis.
Figure 3 illustrates this tradeoff through contour plots of predicted price effects. We compare two scenarios: one where firms achieve maximum cost efficiency (left panel) and another where, besides the maximum efficiency, the merged firm inherits the superior demand characteristics of the two merging firms (right panel). While cost efficiencies can drive price reductions, changes in post-merger conduct often dominate these benefits. Demand-side improvements are relatively minor compared to cost efficiencies and conduct changes. Lastly, we note that the simulated maximum prices of approximately \$229 in Figure 3 align closely with the observed (raw) average prices in treated markets (see Table 2), suggesting the model generates economically meaningful predictions. The contour plots consistently indicate higher post-merger conduct across both efficiency scenarios, though the magnitude of price effects varies with the assumed efficiency gains.
This empirical exercise reveals a fundamental tension in evaluating the three airline mergers. The estimated price stability could reflect either large efficiency gains offset by increased market power or modest but opposing efficiency and competitive conduct changes. This aspect creates a dilemma for merging airlines–claiming significant efficiencies implies accepting evidence of increased market power while arguing for minimal competitive effects undermines the efficiency justification for the merger. A unified approach makes this tradeoff explicit.
@@ 7 Conclusion
Looking ahead, we see promising directions for future research include extending our framework to incorporate endogenous market structure (e.g., Collard-Wexler, 2014; Ciliberto et al., 2021; Fan and Yang, Forthcoming), product characteristics and allowing these to change with mergers (e.g., Sweeting, 2010; Fan, 2013; Wollmann, 2018; Li et al., 2022; Bontemps et al., 2023; Atalay et al., 2024). The challenge will be maintaining our current approach’s transparency while adding these additional features. Our methodology of connecting reduced-form and structural approaches provides a valuable template for such extensions.
@@ A.3 Additional Data Tables
  • Notes: Summary statistics table of the variables (and their units of measurement) used to estimate the structural model in Section 4. For each merger, the sample includes all pairwise combination of the top 50 populated origin and destination markets, restricted to three years leading up to the merger.
@@
  • Notes: Summary statistics table of the variables used to estimate the nested-logit discrete choice model in Section 6, separated by the merger. For each merger, the sample includes all pairwise combination of the top 50 populated origin and destination markets, restricted to three years leading up to the merger.
@@ B.1 DiD Estimation with Levels
  • Notes: This table reports DiD estimates from Equation (1) for airfare, passenger volume and offered seats. These variables are measured in levels, where seats are in 1,000. The analysis is structured such that the first three columns show the estimates for the three mergers separately, and the last column shows the estimates for the “pooled” sample. Panel A presents results using mean-deviation fixed effects, and Panel B presents results using the first-differences method. ${}^{***}:p<0.01$; ${}^{**}:p<0.05$.
The estimates are displayed in Table B.1. Panel A shows substantial merger effects when using the fixed-effects approach. The DL-NW merger reduced fares by \$6.69 ($\approx 3\%=(6.69/225.39)\%$ relative to pre-merger fares) while increasing passengers by 3,728 and capacity by 34,000 seats per market. These effects are both economically and statistically significant. The UA-CO merger showed minimal price effects (\$1.76 increase, statistically insignificant) but similar capacity expansion. The AA-US merger generated the largest fare reduction of \$16.14 (6.8%) with a significant increase in passenger volume and offered seats.
However, these effects largely disappear under the first differences (Panel B) method. The DL-NW fare effect drops to an insignificant -\$0.77, while UA-CO shows a modest -\$1.66 decrease. The AA-US fare effect shrinks to an insignificant -\$1.30 (from a substantial increase of \$32.9 under the fixed effects method). Similarly, passenger and capacity effects attenuate by 80-85%, though remaining statistically significant.
These findings provide evidence of stable post-merger prices and align with our main findings in Section 3 Table 3, indicating more modest merger effects than traditional fixed-effects approaches might suggest.
@@ B.2 Analysis at the Airline-Market Level
For our main analysis presented in Table 3, we estimated Equation (1) at the market level, where each observation represents a unique market-time pair (e.g., the logarithm of average flight prices for Atlanta to St. Louis market at a given time). As a robustness check, we consider the fixed-effects (mean deviation) estimator to examine whether the competitive merger effects documented in Table 3 Panel A might have been driven by market-level aggregation. In particular, we extend our analysis by estimating merger effects at the more granular market-airline level. In this specification, each observation corresponds to a unique market-airline-time combination—for instance, the logarithm of average prices for flights operated by Delta in the ATL-STL market at a specific period.
Our primary variable of interest remains $(\text{Treated}_{mt}\times\text{Post-Merger}_{mt})$ as defined in Equation (1). However, instead of incorporating market-level fixed effects, we include “market-airline” fixed effects while retaining the time fixed effects. Similarly, the % nonstop now measures the percentage of nonstop passengers for a specific airline in a given market.
@@ B.3 Pre- and Post-Merger Windows
  • Notes: This table reports DiD estimates from Equation (1) for airfare, passenger volume and offered seats. Panel A presents results using one month before/after as relevant period to estimate merger effects across, and Panel B presents results using two months before/after as relevant period. ${}^{***}:p<0.01$; ${}^{**}:p<0.05$.
We present the results in Table B.3. Panel A shows that when we use only one month before/after as the relevant period, the merger effects on airfares are imprecisely estimated across all merger cases. The effects on passenger volume appear mixed, with significant negative effects for DL-NW, insignificant effects for UA-CO, and marginally significant positive effects for AA-US. However, capacity (offered seats) consistently shows significant increases in overlap routes across all mergers. Panel B presents results using a two-month window. The pattern is similar—airfare effects remain statistically insignificant, passenger volume effects vary across merger cases, and capacity effects maintain positive significance across all mergers. These findings suggest that the immediate effects of mergers differ from longer-term impacts, with capacity adjustments occurring more rapidly than price and demand responses. The consistent positive effect on offered seats across both time windows indicates that merged airlines quickly increased capacity in overlapping markets. However, the corresponding price adjustments and passenger volumes may require more time to materialize or stabilize.
@@ B.5 Pre-trend Analysis
In Figure B.1, we present the estimates from pre-trend analysis for 2-to-1 markets, mirroring the methodology employed in Figure 2 for 3-to-2 markets. (The UA-CO merger is excluded because there are no 2-to-1 treated markets.) The results reveal that the merger coefficients for price and passenger volume exhibit substantial instability in pre-merger periods across the two mergers. This instability suggests a violation of the parallel trends assumption.
@@ B.6 Synthetic DiD
Recent methodological innovations offer a promising avenue to address this issue. The synthetic control method (Abadie and Gardeazabal, 2003) extended to the DiD framework by Arkhangelsky et al. (2021) provides a data-driven approach to constructing control markets. This method systematically combines potential control markets to create a synthetic comparison group that better approximates the counterfactual of the treated markets. Providing a systematic method to construct control groups reduces researcher discretion in market selection and relaxes the rigid assumptions about trend specifications inherent in our previous analysis. For instance, when one is analyzing the DL-NW merger, a treated market with pre-merger characteristics similar to DL-NW overlap markets would receive a higher weight than a dissimilar market without parametric assumptions about the evolution of market outcomes.
@@
We apply this approach using the implementation package developed by Clarke et al. (2023) to reassess the findings in Table 3. The estimates are presented in Table B.5. These effects suggest three key results. First, we find limited evidence of merger effects on prices and passenger volumes. For instance, the UA-CO merger shows a minimal fare effect of -0.4% in levels and -0.8% in first differences (the latter being marginally significant). The passenger volume effects are similarly modest, approximately 1% in both specifications, and statistically indistinguishable from zero. These estimates suggest economically negligible price impacts from the merger.